OpenStudy (anonymous):

Somebody help please! Add and simplify (3/tw^2)+(4/t^2w)

OpenStudy (saifoo.khan):

\[\frac{3}{tw^2}+\frac{4}{t^2w}\]

OpenStudy (anonymous):

(3t+4w)/(wt)^2

OpenStudy (saifoo.khan):

LCM = t^w^ \[\frac{3t+4w}{t^2w^2}\]